How many grams of solute are in each of the following solutions?

1. 12.5 mL of a 1.50 M AgNO3 solution

2. 24.2 mL of a 7.00 M H3PO4 solution
1
Expert's answer
2020-03-16T12:12:44-0400


  1. molarity = moles per litre.

if 1.50moles of AgNO3 are in 1000ml then 0.01875 moles are in 12.5ml of AgNO3

"mass=RFM \times moles"

mass=170* 0.01875= 3.1875g

2 . if 7.00moles are in 1000ml of H3PO4 solution then 0.1694 moles are in 24.2ml

mass= 98 *0.1694 = 16.6012g